OpenStudy (anonymous):

A certain radioactive isotope has a half-life of approximately 1850 years. How many years to the nearest year would be required for a given amount of this isotope to decay to 40% of that amount?

OpenStudy (anonymous):

N = A (0.5)^(t/j) j = half life period = 1850 years so amount left = 40% of original N = 40% A N = 0.4 A Therefore 0.4 A = A (0.5) ^ (t/1850) 0.5 ^ (t/1850) = 0.4 log (0.5 ^ (t/1850)) = log 0.4 t/1850 = log 0.4 / log 0.5 t = (1850*log0.4) / log 0.5 Solve this,,, to get t = 2445.566 years